OpenStudy (help_people):

A library building is in the shape of a rectangle. Its floor has a length of (3x + 5) meters and a width of (5x − 1) meters. The expression below represents the area of the floor of the building in square meters: (3x + 5)(5x − 1) Which of the following simplified expressions represents the area of the floor of the library building in square meters? 28x − 5 15x2 − 5 15x2 + 28x − 5 15x2 + 22x − 5

OpenStudy (anonymous):

Use the FOIL method 

OpenStudy (anonymous):

(3x+5)(5x−1) =(3x+5)(5x+−1) =(3x)(5x)+(3x)(−1)+(5)(5x)+(5)(−1) =15x2−3x+25x−5 =15x^2+22x−5

OpenStudy (help_people):

THANK YOUUUUUU :) @sunnychick

OpenStudy (anonymous):

No problem my friend :)